OpenStudy (anonymous):

Which logarithmic graph can be used to approximate the value of y in the equation 3y = 8? The graph of function f of x equals log base 4 of x minus 4 The graph of function f of x equals log base 3 of x The graph of function f of x equals log base 4 of x The graph of function f of x equals log base 4 of x plus 3 @StudyGurl14

OpenStudy (anonymous):

\[3^y=8\\ log_3(8)=y\]

OpenStudy (anonymous):

so i guess you are supposed to say \[f(x)=\log_3(x)\]

OpenStudy (anonymous):

then finding \[\log_3(8)\] is the same as solving \(3^y=8\)

OpenStudy (here_to_help15):

you are basically looking for a base 3 log graph

OpenStudy (here_to_help15):

take the base 3 log on both sides of 3^y=8 to get y=base 3 log 8

OpenStudy (here_to_help15):

to find the curve that is base three log 8 you should look at some points; for instance base 3 log 9=2, and base 3 log 3=1

OpenStudy (here_to_help15):

So you should get a graph with points (1,0) (3,1) (9,2)

OpenStudy (here_to_help15):

I can tell you right now that 8 is a lot closer to 9 than it is 3, so you should see a point (8,approx. 2)

OpenStudy (here_to_help15):

Hello?